OpenStudy (anonymous):

Use the Remainder Theorem to evaluate f(x)=x^5+2x^4−2x+10 when x=3. A. 328 B. 43 C. 13 D. 409

OpenStudy (kc_kennylau):

Just plug x inside f(x) :) \[f(\color{green}x)=\color{green}x^5+2\color{green}x^4-2\color{green}x+10\]\[f(\color{green}3)=\color{green}3^5+2(\color{green}3)^4-2(\color{green}3)+10\]\[\hspace{21pt}=243+2\times81-6+10\]\[\hspace{21pt}=\cdots\]

OpenStudy (kc_kennylau):

Do you have problem in the \(\cdots\) part ? :)

OpenStudy (anonymous):

409:)